find the zeros of f(x)=-2x^2+4x+1
Answer by stanbon(75887)

find the zeros of f(x)=-2x^2+4x+1
-----
Quadratic equation::
x = [-4 +- sqrt(16-4*-2*1)]/(-4)
-----
x = [-4+- sqrt(24)]/(-4)
-----
x = [-4+-2sqrt(6)]/(-4)
------
x = 1 -(1/2)sqrt(6) or x = 1 +(1/2)sqrt(6)
--------------
Cheers,
Stan H.
-----------